Trucks in a delivery fleet travel a mean of 100 miles per day with a standard deviation of 29 miles per day. The mileage per day is distributed normally. Find the probability that a truck drives between 120 and 158 miles in a day. Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = = 100

standard deviation = = 29

P(120 < x <158 ) = P((120 - 100) /29 ) < (x - ) /  < (158 - 100) / 29) )

= P(0.69 < z < 2)

= P(z < 2) - P(z < 0.69)

= 0.9772 - 0.7549

= 0.2223

Probability = 0.2223